Systems Staff Meeting Notes 01/05/04

Systems Staff Meeting Notes
January 5, 2004

Present: Rich Bennett, Cynthia Bowker, Will Chaney, Bill Covey, Debra Fetzer, Winston Harris, Bobbie Meng, Hilton Soto, Geoffrey Stewart, Ying Tang
Announcements
  1. Bobbie found gold/diamond teeth at the beach (with her metal detector).
  2. ALA is this week.
  3. Ex-Libris got severely hacked – hit eight customer servers that, in turn, attacked customers. This was all covered up for about 4-6 weeks. It was a major security breach – Ex Libris had very amateurish security in place for the world’s leading library company. Luckily this did not involve us.
Moves/Construction Status
  1. Gary said that we could pull all of the public stuff from West. Discussion on whether there is still Circ activity in West. There will probably be Circ going on there until LAD is ready.
  2. DOT is not ready for the wiring crews yet. Two pipes are in; construction crews are currently working where the remaining 2 pipes need to go. CNS said they would be finished by 12/26, but it was not so. They will have to wait for the construction to terminate before they can finish the wiring.
  3. Books are not scheduled to move to DOT until the end of this month. Circ stuff probably won’t be pulled until the end of February. Details continue to change.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. Barcode Readers – Patrick from Journalism was requesting two – he needs to go through SysRequest. All new ones read GatorOne cards; we no longer purchase ones that don’t. Discussion evolves about increasing abuse of scanners. Bill will make mention to supervisors. The last two scanners were given to Gary. Four more have been ordered. Bobbie has a stack of returned ones that she needs to go through and test, whenever the day comes that she gets a chance.
Liaisons
  1. New additions to the corps.
Networking
  1. NERSP reboot in the middle of the day on Friday.
  2. Server room renovation – network done, but no power or AC.
  3. Patch pusher – working for the most part – Will needs to verify a series of failing over - on the same machines. He has to push those manually. Should be fixed by the next patch push.
  4. Active Directory continues right along.
Software Projects Status
  1. Storage Retrieval – Hot button issue: access through non-IE browsers. Specifically, can’t access it via Mac. Came up with a single request page rather than multiple for non-IE. Tested on Mac, DLC & Netscape – all are fine. Have yet to test Linux. More selection fields were customized on the multiple request.
  2. ID Checker for HSS – now complete. Suddenly their entire staff needed access.
  3. DLC – Maps server stopped working just before Christmas. Joe patched it and had a large data load – after that, it did not work. TomCat worked fine, it was not the cause.
  4. DLC Tracking Database – ongoing.
  5. Mark left Ying a message today regarding converting two Access apps to SQL. He is going to distribute them to other institutions – if so, it is up to him to support.
Web Projects
  1. Rich has been monitoring Web stuff while Tom has been out. That account gets more spam than any other.
DP Coordinator
  1. New student hired: Jessica Patronik. Classes and students (Nicole, Nikia & Jessica) resume tomorrow.


Back to the Systems Homepage

Last updated January 12, 2004
by Debra Fetzer
debharr@mail.uflib.ufl.edu